OpenStudy (anonymous):

100 cm^3 of Nitrogen gas have a mass of 0.14 g. 1) Calculate the molar mass of a gaseous hydrocarbon if 250 cm^3 of this gas weighs 0.7 g at the same temperature and pressure of nitrogen. Thank you! will give medal!

OpenStudy (anonymous):

Well, PV=nRT, and you know that both systems are at the same P and T. Thus, two separate equations can be equated. P/T= (nR)/V Now given that R is a constant and n is the only variable we have (nR)/Vn [ Nitrogen] = (nR)/V [Hydrocarbon] n for nitrogen is known, since the mass number 14 is known. Solve for n Hydrocarbon. Given that you have n and the mass that n corresponds to in g . the molar mass in g/mol should be easy to find Good Luck

OpenStudy (anonymous):

i still don't understand what to do....